OpenStudy (anonymous):

What is the circumference of a circle whose area is 36π?

OpenStudy (haseeb96):

area = 36 pi area of circle is pi*r^2 now put it at the place of area pi*r^2= 36*pi r= 6 as you know the circumference of circle is 2*pi*r put r=6 you get 2*pi*6 12*pi This is your answer